Hey guys..I’m currently preparing for jee main 2017..My strategy is to complete studying all concepts by Feb Mid..So that I can devote entire March for revision and question paper solving..I’m using books like Das Gupta,ML Khanna,SL Loney,HC Verma,DC Pandey, Resnick Haliday Walker and NCERT and P Bahadur objective chemistry..I’m also using Disha publications 36 years question papers..I Just want to know if This strategy is good or should I shift to something effective..Also I would like to point out that  I completed 12 th last year and got a 99 in my JEE mains.
 

Grade:12th pass

7 Answers

Umakant biswal
7 years ago
dear mahesh 
you are in a right direction , proceed on . 
but u need not to follow all of the books , thats even not required to clear advanced , and this can loose most of your time , 
just for physics , h c verma , and disha  prev year paper , 
maths – m l khanna + prev year paper , + s l loney for integral calculus 
chemi – ncert for inorganic , organic take pearson , physical – p bahadur 
and daily give mock test or attend some kind of ques paper , else , lack of revision will lead to intermixing of concept , 
so,revise the content regularly , 
ALL THE BEST 
REGARDS 
Mahesh Mridula
7 years ago
Sir , Can you also Give a planned regular schedule or timetable including timing for the next 1 month for complete revison and learning concepts of 11 th and 12th..I can work to 14-15 hours a day and I dnt have boards too...Please do the needful Sir..
Umakant biswal
7 years ago
dear mahesh 
i can help you with your daily planning , ( i mean when you have to do which subject ) kindly give me your wake up time and sleeping time , and how much hour u want for recretion , 
REGARDS 
Mahesh Mridula
7 years ago
Sir..I start learning by 4 in the morning and finish by 11..On the whole I would require 2 hours of recreation per day..I can sit for a complete stretch of 1 hour 
Umakant biswal
7 years ago
dear mahesh 
in morning u need to start at 4 . 
4;00 – 4;40 = meditation , 
4;45= 5;30 = revision of previous day (maths ) 
5;30 – 6;45 = daily work
7;00 – 8;30 = start a new chapter ( physics ) 
8;45 – 9;50 = revision of previous day ( chemistry ) 
10;05 = 11;10 = start a new chap ( maths ) 
11;10 – 11;45 = cont . with maths 
12;00 – 1;10 = prev day revision (physics) 
1;10 = 2;15 = mid day meal time 
2;30 – 3;45 = start a new chapter ( maths ) 
3;45 – 5;20 = play outdorr games ( recretion ) 
5;30 = 630 = prev year ques practice ( maths ) 
6;45 – 7;50 = give mock test 
8;00 – 9;10 – prev year paper practice ( physics ) 
9;20 = 10;30 = revise all the formula u studied in whole day ( only formula . keep a separate book for that ) 
10;30 = 11;10 = attend a small mock test on chemistry . 
11;10 – SLEEP 
chapter u need to decide acc to weightage , and your completation level 
ALL THE BEST 
REGARDS 
Mahesh Mridula
7 years ago
Sir..I have this book called I.I.T MATHEMATICS by ML KHANNA and JN SHARMA..It’s a very old book and i guess it was revised in the year 1983..Should I carry on with this book or Use Das Gupta ? This book has nearly 1500 pages..
Jessica J
6 years ago
Hi Mahesh,
I hope you followed up with the strategy and tips which were shared. May i know how much did you score in JEE Main 2017.